When a 58-g tennis ball is served, it accelerates from rest to a speed of 45 m/s. The impact with the racket gives t he ball a constant acceleration over a distance of 44 cm. What is the magnitude of the net force acting on the ball?
Answer #1

using v^2 - u^2 = 2ad
45^2 -    0 = 2 x a x 0.44
a = 2301.14 m/s2

F = ma   = 58 x 10-3 x 2301.14 = 133.47 N
Answer #2

a= v^2/(2*s)= 45^2/(2*0.44)

==> F=ma= 0.058*(45^2/(2*0.44))=133.4 N